Aqueous solution of urea is 20% by mass of solution. Then the percrntage by mass of solvent is

Grade:12th pass

2 Answers

Tanvi
6 years ago
Let the mass of the solution be 100gmSo mass of urea is 20% of 100,so mass of urea is 20 gmSo mass of water = 100 – 20 = 80gmSo mass of urea by solvent(water) = 20/80 x 100% = 25%Regards
Tanvi
6 years ago
Let the mass of the solution be 100gmSo mass of urea is 20% of 100,so mass of urea is 20 gmSo mass of water = 100 – 20 = 80gmSo mass of urea by solvent(water) = 20 X 100 ------- 80 = 25%so the answer is 25%